OpenStudy (anonymous):

If the scale is 1 inche/12 feet, what is the actual length for 4 1/2 inches?

hero (hero):

You are given that 1 inch = 12 ft and asked to find the length 4 + 1/2 inches correct?

OpenStudy (anonymous):

The question states "If the scale is 1 inch over 12 feet (fraction) what is the actual length for 4 feet and 1/2 inches?

OpenStudy (anonymous):

Sorry 4 and 1/2 inches, no feet

hero (hero):

I know what the question says. If you have a scale of 1in/12 ft, then every inch in scale is equivalent to 12 ft in length

OpenStudy (anonymous):

So do I turn 4 1/2 inches into 9/2 and solve??

hero (hero):

4 and 1/2 inches is 4 + 1/2 inches because "and" means "plus"

hero (hero):

I was going to give you a walk through on exactly how to solve it.

OpenStudy (anonymous):

That would be great

hero (hero):

Since 1 inch = 12 ft, what happens if we multiply both sides by 4?

OpenStudy (anonymous):

4/48?

hero (hero):

4 in = 48ft

hero (hero):

by scale

hero (hero):

What happens if we take 1 in = 12ft then divide both sides by 2?

OpenStudy (anonymous):

Confused?? 1/6??

hero (hero):

You'd get (1/2)in = 6ft

hero (hero):

So basically by scaling (4 + 1/2)in = (48 + 6) ft

hero (hero):

So all you have to do is add 48 + 6 to get the correct answer in ft

OpenStudy (anonymous):

I'm going to need to sit and work through it again. Still not making much sense, but I do appreciate you trying to help.